Let V = R3 and let W = {(x, y, z)  R3| z = x + y}. Prove that W is a subspace of V.


1
Expert's answer
2021-11-15T16:28:19-0500


Let V is a linear space and "W\subset V" is some subset of V. W is a linear subspace in V iff:

  1. "\forall(v_1,v_2\in W)(v_1+v_2\in VW;"
  2. "\forall(c\in R,v\in W)(c\cdot v\in W);"

We must verify these two properties.

  1. Let "v_1=(x_1,y_1,z_1) ,v_2=(x_2,y_2,z_2)\in W" . It is means that "z_1=x_1+y_1, \space z_2=x_2+y_2" . Then "v_1+v_2=(x_1+x_2,y_1+y_2,z_1+z_2)\in V=R^3" be definition of "R^3" and because of "z_1+z_2=(x_1+y_1)+(x_2+y_2)=(x_1+x_2)+(y_1+y_2)" we have that "v_1+v_2\in W" , first condition takes place.
  2. Let "v=(x,y,z) \in W, c\in R" . It is means that z=x+y. Value "c\cdot v=c\cdot(x,y,z)=(cx,cy,cz)\in V=R^3"

by definition? to verify that "cv\in W" we see that "c\cdot z=c\cdot(x+y)=c\cdot x+c\cdot y" and this means that the second property is proved also.

Thus we have proved that W is a linear subspace.